### v2

# CodeSSM: Towards State Space Models for Code Understanding

Abstract
Although transformers are widely used for various code-specific tasks, they have some significant limitations. In this paper, we investigate State Space Models (SSMs) as a potential alternative to transformers for code understanding tasks, such as code retrieval, classification, and clone detection. Previous research has already demonstrated that SSMs are more compute-efficient than transformers. In our work, we show that SSMs are also more sample-efficient and can effectively extrapolate to longer contexts (beyond the pretraining context) during fine-tuning. Through comprehensive experiments, we demonstrate that SSMs could serve as a viable alternative to transformers for code understanding tasks, while addressing some of the major limitations associated with transformers.
CodeSSM: Towards State Space Models for Code Understanding
Shweta Verma1, * , Abhinav Anand1, * , Mira Mezini1,2,3 1TU Darmstadt, 2Hessian Center for Artificial Intelligence, Darmstadt, Germany, 3National Research Center for Applied Cybersecurity ATHENE *Equal Contribution, Correspondence: shweta.verma@tu-darmstadt.de
1 Introduction
Transformers (Vaswani et al., 2017) have been known to perform well on various code intelligence tasks. One of the main reasons for this improvement is the pretraining - finetuning paradigm used to train transformer models, which is possible due to the ease of parallelization of attention-based transformer architectures. Under this paradigm, the transformer model is initially trained on a large corpus of unlabeled data using a self-supervised training objective. This is followed by a supervised finetuning process on a smaller labeled dataset.
However, the performance gain offered by the transformer models comes with some trade-offs, such as quadratic complexity, a substantial data requirement, and high inference costs. While methods such as linear attention (Katharopoulos et al., 2020) and sparse attention (Guo et al., 2022b; Zaheer et al., 2020; Condevaux and Harispe, 2022) have been proposed to address computational inefficiency, the practical gains from these methods are limited (Yang et al., 2025; Qin et al., 2022).
Another limitation of transformers is their fixed context window resulting from positional embedding. Although methods such as AliBi (Press et al., 2022) and RoPE (Su et al., 2024) solve the problem of a fixed context window, transformer models still fail to generalize to lengths not seen during pretraining (Peng et al., 2024; Goel et al., 2025). To overcome this challenge, additional techniques are required to extend the context window without significantly compromising performance (Chen et al., 2023b; Hua et al., 2025a; Lin et al., 2024).
One-dimensional Convolutional Neural Networks (1D CNNs) are a fast and positional embedding-free alternative to transformers. 1D CNNs, when trained using the pretraining-finetuning approach, show improved performance on some tasks but fail to capture long-range dependencies in input data (Tay et al., 2021).
State-space models (SSMs)Gu et al. (2021, 2022b, 2022a) represent an alternative that addresses the limitations of transformers and 1D CNNs. SSMs leverage the strengths of 1D CNNs while effectively capturing long-range dependencies with linear-time complexity. Additionally, SSMs are position-aware, eliminating the need for positional embedding.
In this work, we explore the application of SSMs to the domain of code understanding. To achieve this, we pretrain an encoder-only model based on SSM, which we call CodeSSM, on a small code dataset with masked language modeling (MLM) (Devlin et al., 2019). We evaluated our model on multiple code understanding benchmarks. We also investigate various training aspects, such as the effects of positional embedding, dropout, bidirectional SSMs, and varying pretraining context lengths. We compare the CodeSSM with two transformer models that we train – (1) BertCoder: based on Bert architecture (Devlin et al., 2019) and (2) RoCoder: based on RoFormer architecture Su et al. (2024). We assess the performance of our model against several high-performing transformers as well.
Our study demonstrates that SSMs exhibit strong sample efficiency in code language modeling tasks. Notably, CodeSSM achieves over 50% accuracy in masked language modeling (MLM) with fewer than 3,000 training steps.
Additionally, the code language modeling capabilities of CodeSSM also transfer well to downstream tasks as CodeSSM outperforms transformers on multiple code understanding tasks when pretrained under similar conditions. CodeSSM is also competitive with (and even surpasses) transformer models trained on significantly larger datasets and with more complex pretraining objectives. Furthermore, another advantage of SSMs is their ability to extrapolate to much larger contexts than those encountered during pretraining, effectively addressing a key limitation of the transformer architecture.
Our objective in training the CodeSSM model is not to achieve state-of-the-art performance on coding benchmarks, but to explore and document the strengths and limitations of SSMs in understanding code. This goal also motivates our usage of a small dataset and MLM pretraining which allows us to quickly experiment with different model configurations and compare them under similar training settings. Although we observe improvements in performance with larger pretraining data, we will leave the exploration of very large training datasets and more complex training objectives for future work.
Summary of contributions:
  • We conduct the first systematic evaluation of State Space Models (SSMs) for code understanding, introducing a specialized model, which we term CodeSSM. Our analysis reveals that CodeSSM achieves robust performance across multiple benchmarks, consistently outperforming attention-based models while requiring substantially less pretraining data.
  • Our study presents the first empirical evidence that State Space Models (SSMs) achieve superior sample efficiency in code-understanding tasks compared to transformers. CodeSSM substantially outperforms transformer baselines in data-constrained scenarios with only a few thousand training samples.
  • We demonstrate that CodeSSM addresses a key limitation of transformer architectures –extrapolation to unseen context lengths. By examining the effects of different pretraining context lengths and positional embeddings in CodeSSM, we show that SSMs outperform transformers as downstream input sequences become longer.
  • We observe that including even a small dropout mask in the CodeSSM layer degrades the performance of the model even on very small benchmarks. CodeSSM also performs very well on the adversarial text-code search task. Together these observations indicate that SSMs exhibit robust semantic understanding capabilities.
2 CodeSSM
Figure 1: CodeSSM layer. Transform represents SSMs in CodeSSMs and Discrete Fourier Transforms in CodeF. Dropout is only present in one of the CodeSSM variants, CodeSSM-do. In CodeSSM-uni, both the flip operations are removed. Non-linearity is not shown for clarity.
In this section, we introduce the basic architecture of CodeSSM and its variations that we investigate. We also explain the pretraining setup.
2.1 Architecture
CodeSSM is an encoder-only model consisting of 12 layers  [^1] The trained weights are available here: https://figshare.com/s/14238287e9078f92cd50.. This model is built upon the Bidirectional Gated SSM (BiGS) architecture Wang et al. (2023a), incorporating certain variations that we examine and their applicability in the field of code understanding.
CodeSSM-pos. Transformer models Vaswani et al. (2017) require positional embeddings to address their permutation invariance in self-attention layers. Since SSMs are inherently position-aware, they do not require positional embeddings. However, to understand the implications of positional embeddings on downstream performance, we include them in one variant of CodeSSM, which we call CodeSSM-pos.
CodeSSM-dropout. Dropout layers (Srivastava et al., 2014) prevent over-fitting in neural networks. So, to investigate over-fitting in CodeSSM, we create a variant with two dropout masks, positioned after each multiplication operations (see Figure 1). We refer to this variant as CodeSSM-do.
CodeF and CodeSSM-Uni. Recent research Lee-Thorp et al. (2022) questioned the necessity of self-attention in transformers. To investigate the significance of SSM, we conducted a similar ablation study by replacing SSM blocks in CodeSSM with DFT (see Figure 1), creating a variant we call CodeF. Furthermore, we analyze the importance of bidirectional processing via CodeSSM-Uni (Unidirectional Gated SSM), a variant of CodeSSM that removes flip operations from the CodeSSM layer.
2.2 Pretraining
We trained all models with the MLM objective using the MLM setup proposed by Devlin et al. (2019). We use different pretraining setups to investigate various aspects of SSMs on code understanding tasks. Here, we explain this setup.
Dataset. We initially trained CodeSSM and a BERT-like transformer model, which we refer to as BertCoder, on 1.8M text-code pairs from the CodeSearchNet (CSN) dataset. The small pretraining dataset allows us to quickly experiment with various model configurations. Both models were trained with a large context window of 2048, enabling them to be fine-tuned on benchmarks with lengths shorter than 2048. We refer to the CodeSSM model trained with the CSN dataset as CodeSSM-base. Additionally, we also train CodeSSM variants such as CodeSSM-pos, CodeSSM-do, CodeF and CodeSSM-Uni (see 2.1) on the same dataset and context length.
Context Length and Length Extrapolation. To investigate the length extrapolation capability of SSMs, we pretrain the CodeSSM model with a 256-length context window. Since we cannot change the context window for BertCoder once pretrained with a smaller context, we pretrained a RoFormer model Su et al. (2024), which is similar to the BERT model but with Rotatory Positional Embeddings (RoPE), with a 256-length context window. We refer to this model as RoCoder.
CSN dataset separates docstring and code, which creates an issue when they are concatenated and subsequently truncated. This can result in one modality (text) being much longer than the other modality (code). We also observed this discrepancy results in reduction in performance in tasks involving two sequences (refer to column 3 of Table 2).
To alleviate this issue, we use a subset of 1.8M samples from the StarCoder dataset (Li et al., 2023) to train the models with 256-length context (referred to as CodeSSM-sc-256). This subset contains 300k samples from the six programming languages (PLs) in the CSN dataset  [^2] Go, Java, Javascript, PHP, Python, Ruby. We also train CodeSSM on this dataset with a context length of 2048 (referred to as CodeSSM-sc-2048).
Data Scaling. To study the impact of large datasets on the downstream performance of CodeSSM, we also train it on a large subset of the StarCoder dataset with 4.8M samples (referred as CodeSSM-sc-L). This subset contains 600k samples from 8 PLs (C and C# along with CSN PLs).
Table 1: Comparison of dataset size and pretraining objective of various encoder-only models. Replaced Token Detection (RTD), Unidirectional Language Modeling (ULM), Denoising objective (DNS), Data Flow Edge Prediction (DFEP), Node Alignment (NA), Masked Span Prediction (MSP), Identifier Tagging (IP), Masked Identifier Prediction (MIP)
Models Dataset Pretraining
Size Objectives
CodeSSM-base 1.8M CSN MLM
CodeSSM-SC 1.8M SC data MLM
CodeSSM-SC-L 4.8M SC data MLM
CodeBERT 8.5M CSN MLM, RTD
UniXCoder 6.4M CSN + C4 data MLM, ULM, DNS
GraphCodeBERT 6.4M CSN MLM, DFEP
CodeT5 8.35M CSN + C, CSharp MSP, IP, MIP
3 Experiments
In this section, we describe the benchmarks on which we evaluate CodeSSM and the baselines with which we compare our model.
3.1 Baselines
We trained two transformer-based models as baselines: a large context (2048) BERT model to compare SSMs’ large context abilities and a small context RoFormer model to compare the length extrapolation abilities of transformers and SSMs. The models being compared are trained under the same settings: same dataset, same objective, context length, and batch size.
Additionally, we also mention the performance of some well-performing transformer models in each task to place CodeSSM in the context of SOTA encoder-only transformer models. However, it should be noted that these transformer models are trained on a much larger dataset and with a mixture of multiple training objectives (see Table 1).
3.2 Benchmarks
We benchmark the models on seven code understanding tasks: two retrieval tasks, three sequence classification tasks, one token-level classification task, and one clone detection task.
Retrieval tasks. There are two retrieval tasks on which we evaluate our models, NLCodeSearch (AdvTest) Lu et al. (2021) and Stackoverflow Question Answer (SQA). In NLCodeSearch, a description in natural language is provided, and the goal is to find the source code that corresponds to that description. To evaluate a model’s ability to generalize, function names and variable names in the test sets are replaced with generic tokens, for example, the function name with func. This task is difficult because the model has to locate the correct code throughout the test set. In SQA, a StackOverflow question is given and the goal is to retrieve the highest upvoted answer. This dataset is derived from the original StackOverflow dataset  [^3] https://www.kaggle.com/datasets/stackoverflow/stacksample/data. The average length of question and answer are 1.2k and 1.4k respectively.
Sequence classification tasks. Sequence classification includes two vulnerability detection tasks, Devign Zhou et al. (2019) and DiverseVul Chen et al. (2023a), as well as a complexity prediction task Jeon et al. (2023). Vulnerability detection is a binary classification task where the goal is to identify whether a given code contains security vulnerabilities. The Devign benchmark Zhou et al. (2019) is a balanced dataset, while DiverseVul consists of only around 5% vulnerable samples. So, we use accuracy for the former and F1 macro for the latter. Complexity prediction Jeon et al. (2023) is a small benchmark with only 3613 training samples to predict the algorithmic complexity of a given Java code among 7 different labels.
Token classification task. In the type-inference Jesse and Devanbu (2022) task, the model has to predict the type (such as integer, boolean, etc.) of each token in a given code. The label consists of 50K most frequent (built-in and user-defined) types, while the remaining types have been replaced with the UNK type. During the evaluation, predicting the UNK token is considered incorrect.
Clone detection task. In the clone detection task Wang et al. (2020), two pieces of code are given as input, and the task is to perform binary classification to determine whether they are semantically equivalent. This task is challenging due to the highly unbalanced dataset, which contains a significantly high number of semantically dissimilar samples. The models are evaluated using the F1 score.
For all tasks, we maintain the same context length as specified in the original benchmarks. All models are fine-tuned using a context length of 512 for text-to-code search, clone detection, type inference, and DiverseVul. For the Devign benchmark, the models are trained with a context length of 400. In SQA and complexity prediction, we use a context length of 1024.
Table 2: Results on NLCodeSearch (AdvTest), Clone detection and Type inference. The result also includes details about the pretraining dataset and pretraining context length. The best performance is in bold and other noteworthy results are underlined.
Dataset Model Context Length NLCodeSearch Clone Detection Type Inference
(Pretraining) (Pretraining) (MRR) (F1) (Overall F1)
CodeBERT 512 27.19 0.941 0.595
BertCoder 2048 24.30 0.942 0.666
CodeSeachNet CodeSSM-base 256 21.87 0.908 0.588
CodeSSM-base 2048 28.38 0.935 0.615
CodeSSM-pos 2048 26.58 0.939 0.594
CodeSSM-do 2048 26.33 0.931 0.609
RoCoder 256 29.44 0.938 0.687
StarCoder CodeSSM-Sc 256 28.41 0.948 0.624
CodeSSM-Sc 2048 29.88 0.938 0.625
CodeSSM-Sc-L 2048 30.87 0.940 0.637
Table 3: Results on SQA and APPS benchmark in terms of MRR. The best performance is in bold and other noteworthy results are underlined.
Model SQA
BertCoder 76.25
RoCoder 80.91
CodeSSM-Sc-256 82.12
CodeSSM-Sc-2048 82.29
CodeSSM-Sc-L-2048 82.59
4 Results
We present the performance of our model on various benchmarks described in Section 3.2
4.1 Retrieval Tasks
NLCodeSearch(AdvTest). Column 3 of Table 2 presents results of various models evaluated on NLCodeSearch benchmark. Among the models trained on the CSN dataset, CodeSSM-base (with a context size of 2048) outperforms other models, including CodeBERT, which was pretrained with significantly more data. CodeSSM-base with a context length of 256 performs poorly compared to other variants – we explain the reason for this in Section 2.2. CodeSSM-sc-256 shows remarkable performance and length extrapolation to double the length on which it was trained and is competitive with RoCoder and CodeSSM-sc-2048.
These results are particularly noteworthy given the adversarial nature of the NLCodeSearch benchmark, where critical semantic identifiers such as function and variable names are obfuscated with generic placeholder tokens. The superior performance of CodeSSM under these challenging conditions underscores its efficiency in capturing the underlying structure of code, independent of specific identifier names, even if it is pretrained with much less data compared to CodeBERT.
StackOverflowQA. Table 3 presents the performance of CodeSSM in conjunction with the baseline models trained by us, BertCoder and RoCoder, on the SQA benchmark. SQA allows us to evaluate the models with longer contexts. CodeSSM-sc with a pretraining context length of 256 achieves strong results – it successfully extrapolates to contexts four times longer than those seen during pretraining. Furthermore, CodeSSM-sc outperforms BertCoder (pretrained with a context length of 2048) by 6 points, despite being pretrained with a context length eight times shorter. CodeSSM also surpasses RoCoder, highlighting its superior ability to extrapolate to longer contexts and capture long-range dependencies. Furthermore, we observe that increasing the pretraining context length leads to further improvements in CodeSSM’s performance.
Table 4: Results on DiverseVul benchmark in terms of F1 macro. The best performance is in bold and other noteworthy results are underlined.
Model DiverseVul Devign
(F1 macro) (Accuracy)
CodeBERT 37.85 62.08
GraphCodeBERT 36.79 63.21
CodeT5 45.69 62.88
BertCoder 65.97 62.50
RoCoder 68.15 53.47
CodeSSM-Sc-256 69.31 64.53
CodeSSM-Sc-2048 67.72 63.14
CodeSSM-Sc-L-2048 68.76 63.54
Table 5: Results on Complexity Prediction. The best performance is in bold and other noteworthy results are underlined.
Model F1
CodeBERT 85.81
GraphCodeBERT 87.98
UnixCoder 93.75
BertCoder 89.66
RoCoder 93.06
CodeSSM-do-2048 92.82
CodeSSM-Sc-256 93.77
CodeSSM-Sc-2048 94.37
CodeSSM-Sc-L-2048 94.55
4.2 Sequence classification tasks
Devign. Column 3 of Table 4 shows the results of the various models on the devign benchmark. CodeSSM outperforms BertCoder and RoCoder as well as CodeBERT, GraphCodeBERT, and CodeT5. Interestingly, BertCoder, pretrained with a longer context, also outperforms CodeBERT, showcasing the importance of a longer context for transformers. However, in the case of CodeSSM, the smaller context model performs much better. Notably, RoCoder performs poorly, potentially due to over-fitting, as the devign dataset is very small (only 21k training samples). RoPE, used in RoCoder, is also prone to positional overfitting Liu et al. (2024b); Goel et al. (2025).
DiverseVul. Column 2 of Table 4 shows that CodeSSM and its variants outperform BertCoder and RoCoder on the DiverseVul benchmark, which is a larger vulnerability detection dataset compared to Devign. It shows that CodeSSM can perform better than transformers even on a larger dataset. CodeSSM even outperforms CodeBERT, GraphCodeBERT, and CodeT5, which are trained on a much larger dataset and have multiple pretraining objectives (see Table 1).
Complexity Prediction. Table 5 shows that CodeSSM-sc-256 outperforms all the transformer models even the ones trained with longer context length (BertCoder) and larger data and more pretraining objectives (CodeBERT, GraphCodeBERT, UnixCoder). Additionally, CodeSSM-sc-2048 shows very strong performance as pretraining on longer context length improves the performance further on this task.
4.3 Token classification task (Type inference)
Column 5 of Table 2 shows the results on the type inference benchmark. CodeSSM (across all context lengths) matches or exceeds CodeBERT. However, both RoCoder and BertCoder outperform CodeSSM. We provide reasons for poor performance of CodeSSM on type inference in Section 5.5 and Appendix G.
4.4 Clone Detection
Column 4 of Table 2) shows the F1 on clone detection task. The CodeSSM-sc with a pretraining context length of 256 outperforms both CodeBERT, BertCoder, and RoCoder in terms of F1. Under a similar training setup (2048 context on CSN), removing positional embedding does have a minimal impact on performance, but with the advantage that the maximum input length need not be fixed during pretraining. The decrease in performance after introducing dropout shows the model does not overfit despite the dataset being unbalanced.
5 Research Questions
In this section, we discuss some important research questions related to CodeSSM and its characteristics.
5.1 How sample efficient is CodeSSM?
Figure 2: MLM training accuracy of CodeSSM and BertCoder, trained with a context window of 2048. The CodeSSM model achieves decent MLM accuracy in very few training steps.
Figure 2 shows the MLM training accuracy trajectories of CodeSSM and BertCoder, both trained with a maximum context of 2048. CodeSSM demonstrates rapid initial learning, achieving more than 50% accuracy in only 3,000 training steps (96,000 training samples). BertCoder, on the other hand, exhibits better long-term learning characteristics, outperforming CodeSSM by 4% points after six training epochs (see Appendix E). However, CodeSSM outperforms transformers in all downstream tasks except type inference, as discussed in Section 4. The performance on downstream tasks shows that the code modeling capability observed during pretraining transfers well to multiple downstream tasks.
We evaluated CodeSSM on two small benchmarks: The Devign vulnerability detection benchmark with only 21k training samples and the complexity prediction benchmark with only 3.6k training samples. In both benchmarks, CodeSSM not only outperforms BertCoder and RoCoder but also many other models trained with significantly more data and multiple code-based training objectives.
Our analyses demonstrate that SSMs represent a sample-efficient alternative to transformers for code language modeling and understanding tasks. We attribute the sample efficiency of CodeSSM to the distinct inductive bias of SSM models compared to transformer architectures Wang et al. (2023a). We also present some (informal) theoretical basis for CodeSSM’s performance and the magnitude and phase spectrum of Fourier Transform of SSM kernel in Appendix F.
5.2 How well does CodeSSM extrapolate to longer contexts?
We studied the transformer models and CodeSSMs for positional embedding and different pretraining context lengths to investigate their length extrapolation capabilities.
On multiple tasks, we observe that BertCoder outperforms CodeBERT despite less pretraining data simply due to the longer pretraining context length. However, replacing absolute positional embedding with RoPE can improve performance even with a pretraining context of only 256. But even with RoPE, RoCoder fails to match the performance of CodeSSM except in type inference and NLCodeSearch.
Out of two retrieval tasks, RoCoder performs slightly better on NLCodeSearch while CodeSSM outperforms it on SQA. The context length is 512 for the former and 1024 for the latter. Thus, CodeSSM is better at handling unseen input lengths as the context for the downstream task increases. Since SSMs have linear complexity with respect to input length, they are advantageous in terms of both computational efficiency and performance. Moreover, with a longer pretraining context, CodeSSM outperforms RoCoder on NLCodeSearch as well. Therefore, with more computational resources, CodeSSM can be efficiently trained with longer contexts to enhance the performance on retrieval tasks. However, it’s important to note that longer pretraining contexts do not benefit all downstream tasks; in fact, performance slightly decreases on vulnerability detection and clone detection.
Furthermore, we observe that CodeSSM performs better without positional embedding. CodeSSMs without positional embeddings do not face any length limitation, unlike transformers with absolute positional embedding, or positional biases observed with different methods such as Alibi or RoPE (Liu et al., 2024b; Goel et al., 2025).
5.3 Can SSMs be replaced by DFT?
We ask this question because Lee-Thorp et al. (2022) showed that self-attention can be replaced with DFT in transformers with minimal reduction on NLP tasks. We also observed that in all tasks RoCoder performed significantly better than BertCoder. Hua et al. (2025b) showed that RoPE also performs token-level (non-uniform) DFT. Additionally, SSMs also implicitly perform DFT of input and kernel during convolution operation.
To answer this question, we performed an ablation study by replacing the SSM transformation in CodeSSM with DFT and creating a unidirectional variant of CodeSSM. The results are presented in Figure 3 (and in Table 7). We observe that both variants perform significantly worse compared to CodeSSM. However, it is important to note that CodeF performs better than CodeSSM-uni, i.e., the advantages of SSM over DFT on code modeling are only realized with bidirectional SSMs.
Figure 3: MLM training accuracy for different variations to the SSM layer: CodeSSM (Bi-directional), CodeSSM-Uni (Uni-directional) and CodeF (Discrete Fourier Transform in place of SSM). Replacing SSM with DFT reduces the modeling capacity significantly.
5.4 Is CodeSSM robust to overfitting on small datasets?
We study the robustness of the CodeSSM model in two ways. First, we study the impact of adding dropout to the CodeSSM layer. Across all tasks, including the complexity prediction task which has only 3.6k samples, the performance of the model reduces when dropout is added. The performance degradation suggests that CodeSSM does not overfit during training regardless of the size of the finetuning dataset. Second, we evaluated the model on the adversarial NLCodeSearch task, by which informative tokens such as function and variable names have been replaced with noninformative generic tokens. The CodeSSM model performs very well in this task even with minimal training.
The reduction in performance with dropout and the performance of CodeSSM in the adversarial text-to-code search task demonstrate its robust semantic understanding capabilities on code understanding tasks.
5.5 When does CodeSSM fail and why?
We discuss two possible scenarios where the performance of SSMs on code understanding tasks can be inferior to that of the transformers.
As discussed in Section 5.1, transformers can outperform SSMs as the number of samples increases. This phenomenon can hinder the performance of SSMs on some tasks such as Type Inference. This benchmark consists of over 9M type annotations, which is significantly larger than the samples of other tasks. The total number of predictions (hence, feedback to the model through loss) is even larger since there is one prediction for every token. However, when the number of samples is large and the context length of finetuning datasets is long, SSMs can surpass transformers in such scenarios (as seen in the SQA benchmark).
SSMs can effectively capture long-range dependencies but do not model local dependencies (Zuo et al., 2024) very well. The type inference benchmark requires the model to understand both short and long-range dependencies. For example, in the code sample x = y+1; z = func(x), the type of x can be inferred from the local context (variable y), while the type of z can only be inferred from the return type of the function func, which requires capturing long-range dependencies. We also observe that this problem becomes more challenging when the local context is complicated. We provide a qualitative example of CodeSSM’s failure in the task of type inference in Appendix G. We can potentially improve CodeSSM’s performance on this task by making the step size of SSMs (see Appendix A) data dependent Qi et al. (2024) or with multiple SSM kernels (Appendix G), which can improve its understanding of short-range dependencies.
6 Related Work
Transformer models have been used in code intelligence tasks Feng et al. (2020); Guo et al. (2022a); Wang et al. (2021, 2023b). However, they struggle to scale with long sequence lengths due to the quadratic complexity of self-attention Vaswani et al. (2017). Various approaches have been proposed to address quadratic complexity, including linear attention Katharopoulos et al. (2020) and methods using sparse attention patterns Guo et al. (2022b); Zaheer et al. (2020); Condevaux and Harispe (2022). LongCoder Guo et al. (2023) uses the sparse attention mechanism. However, efficient transformers have seen limited practical adoption due to a minimal gain in computational cost in practice Yang et al. (2025) and a potential degradation in performance Qin et al. (2022). Dao et al. (2022; 2024) introduced hardware-level optimizations, but the benefits of these optimizations are hardware-specific and do not necessarily apply to newer or different hardware architectures Shah et al. (2024).
Moreover, self-attention models are trained with a fixed context window and cannot be used beyond the pretraining context window. Techniques such as ALiBi Press et al. (2022) and RoPE Su et al. (2024) have been proposed to alleviate this problem. HiRoPE Zhang et al. (2024) adapted RoPE to code models. However, these methods come with their own set of challenges Chen et al. (2023b); Hua et al. (2025a); Lin et al. (2024).
In contrast, SSMs, and hence CodeSSM, are efficient by design and have been shown to better capture long-range dependencies compared to self-attention Gu et al. (2022b). Additionally, CodeSSM does not have positional embedding, alleviating the issues of position-specific biases and length generalization.
Guan et al. (2024) and Wang et al. (2024) enhances the base transformer models for longer contexts using repository-level information retrieval, while Clement et al. (2021); Cheng et al. (2024); Lomshakov et al. (2024); Bi et al. (2024); Liu et al. (2024a); Ma et al. (2024) use static analysis tools. The usage of static analysis tools is complementary to CodeSSM, as the performance of CodeSSM can also be further improved through the use of these tools. Additionally, CodeSSM can be a drop-in replacement as an efficient retriever model.
7 Conclusions
State-space models (SSMs) have emerged as an efficient alternative to transformers. In this work, we investigated the effectiveness of SSM in code understanding tasks by introducing CodeSSM. Our results show significant potential of CodeSSM in code modeling and understanding. Our findings show that SSMs can be a robust and sample-efficient alternative to transformers, offering the added benefit of length extrapolation.
While the training pipeline and architectural improvements for transformers have evolved extensively over the past decade, the optimal training conditions for SSMs have yet to be thoroughly investigated. This work represents a step toward understanding these conditions and uncovering the advantages of SSMs. Notably, the performance gains achieved by CodeSSM were obtained under suboptimal conditions, indicating the potential for further enhancements. With advancements in training methodologies and architectural designs for SSM-based models, their performance could be significantly improved.
Limitations
While we have done an extensive study of SSMs for code understanding, our work has some limitations.
  • Single kernel transformation: We have used a simple implementation of S4d Gu et al. (2022a) in the CodeSSM layer. This implementation has only one SSM kernel shared across all input channel. Incorporating multiple kernels in an SSM block can further improve the performance.
  • Short-range dependencies: We have discussed SSMs’ limitations in understanding short-range dependencies with respect to type inference task. However, we do not implement a solution to this problem. Some methods to alleviate this issue have been proposed in literature. We leave the exploration of these methods in the context of code understanding for future work.
  • Generative tasks: CodeSSM currently uses SSMs only in convolution mode. Due to bi-directional SSMs, it cannot be used for auto-regressive tasks. As a result, it is not applicable to generative tasks. Hence, we cannot compare CodeSSM with Mamba Gu and Dao (2024) or Codestral Mamba  [^4] https://mistral.ai/news/codestral-mamba. To alleviate this limitation, future work can utilize the recurrent view of SSM to enable generative capabilities.
Ethical Considerations
We have used publicly existing datasets. CodeSearchNet dataset can contain personally identifiable information (PII). While StarCoder dataset attempts to remove PII, there might still be some PII in the dataset. It is not clear whether our model can leak such data if present in the pre-training dataset, but we acknowledge that such potential risk exist. Further, our work involved training multiple models. While the model and dataset sizes were small, the training can still have environmental impacts.
Appendix A Background
A.1 State Space Model
A state space model (SSM for short) maps a 1-dimensional input signal $u(t)$ to an N-dimensional latent state $x(t)$ and then projects it back to a 1-dimensional output. The transformations follow the equations
$\begin{split}x^{\prime}(t)&=\mathbf{A}x(t)+\mathbf{B}u(t)\\ y(t)&=\mathbf{C}x(t)+\mathbf{D}u(t)\end{split}$ (1)
where $\mathbf{A}$ controls how the system evolves, and $\mathbf{B}$, $\mathbf{C}$, $\mathbf{D}$ are projection parameters. Previous work Gu et al. (2021, 2022b, 2022a) considered SSM as a black-box representation within a deep sequence model. In this context, the parameters $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$, $\mathbf{D}$ are learned with gradient descent. For simplicity, we can assume $\mathbf{D}=0$, as it corresponds to a skip connection and is straightforward to compute. To apply these models to discrete input sequences, Eq. 1 is discretized with a step size of $\Delta$.
$\begin{split}x_{k}&=\overline{\mathbf{A}}x_{k-1}+\overline{\mathbf{B}}u_{k}\\ y_{k}&=\overline{\mathbf{C}}x_{k}\end{split}$ (2)
Eq. 2 can be interpreted as a convolution between the input sequence and a convolution kernel, which is a function of $\mathbf{A}$, $\mathbf{B}$ and $\mathbf{C}$. This convolution view of SSM is expressed as follows :
$\begin{split}y_{k}&=\overline{\mathbf{C}}\overline{\mathbf{A}}^{k}\overline{\mathbf{B}}u_{0}+\overline{\mathbf{C}}\overline{\mathbf{A}}^{k-1}\overline{\mathbf{B}}u_{1}+...+\\ &\overline{\mathbf{C}}\overline{\mathbf{A}}\overline{\mathbf{B}}u_{k-1}+\overline{\mathbf{C}}\overline{\mathbf{B}}u_{k}\\ y&=\overline{\mathbf{K}}*u\\ \overline{\mathbf{K}}&\in\mathbb{R}\coloneqq K_{L}(\overline{\mathbf{A}},\overline{\mathbf{B}},\overline{\mathbf{C}})\end{split}$ (3)
where $\overline{\mathbf{K}}$ is the SSM convolution kernel.
Throughout the paper, we denote the input length as $L$ and the hidden dimension as $d$. We follow Gu et al. (2022a) for all the parameterization.
A.2 BiGS Model
BiGS model Wang et al. (2023a) is a bidirectional gated architecture which consists of two SSM layers as shown in Fig. 1. This gated architecture is adapted from the gated attention unit described in Hua et al. (2022).
A single layer of the BiGS model has a three stage computation. The first stage utilizes a feed-forward layer alongside gating, employing the GELU activation function Hendrycks and Gimpel (2016). The second stage consists of two sequential SSM blocks with a multiplicative gating. The third stage again uses feed-forward and gating.
$\begin{split}\mathbf{X}&=\text{LayerNorm}(\mathbf{X}_{i})\quad\in\mathbb{R}^{L\times d}\\ \mathbf{V}&=\sigma(\mathbf{W}_{v}\mathbf{X})\quad\in\mathbb{R}^{L\times 3d}\\ \mathbf{F}&=\sigma(\mathbf{W}_{f}\mathbf{X})\quad\in\mathbb{R}^{L\times d}\\ \mathbf{B}&=\sigma(\mathbf{W}_{b}\text{Flip}(\mathbf{X}))\quad\in\mathbb{R}^{L\times d}\end{split}$ (4)
$\begin{split}\mathbf{U}_{1}&=\mathbf{W}_{u_{1}}\text{SSM}(\mathbf{F})\quad\in\mathbb{R}^{L\times d}\\ \mathbf{U}_{2}&=\mathbf{W}_{u_{2}}\text{SSM}(\mathbf{B})\quad\in\mathbb{R}^{L\times d}\\ \mathbf{U}&=\sigma(\mathbf{W}_{u}(\mathbf{U}_{1}\otimes\text{Flip}(\mathbf{U}_{2})))\quad\in\mathbb{R}^{L\times 3d}\end{split}$ (5)
$\begin{split}\mathbf{O}&=\mathbf{W}_{o}(\mathbf{U}\otimes\mathbf{V})\quad\in\mathbb{R}^{L\times d}\\ \mathbf{X}_{i+1}&=\mathbf{O}+\mathbf{X}_{i}\quad\in\mathbb{R}^{L\times d}\end{split}$ (6)
The BiGS model consists of 23 layers and includes a positional embedding.
A.3 Model Size
All the models we used, CodeSSM, BertCoder and RoCoder, has 190M parameters.
Appendix B Hardware Details
We run all our experiments on Nvidia A100 GPUs with 80GB memory. All pretraining is done on 4 GPUs and all finetuning is done on a single GPU.
Appendix C Dataset
CodeSearchNet Dataset. The CodeSearchNet dataset (Husain et al., 2019) used in this work is the unofficial dataset available on Huggingface hub  [^5] https://huggingface.co/datasets/code-search-net/code_search_net using datasets library  [^6] https://github.com/huggingface/datasets, since the official dataset is no longer accessible. The unofficial dataset is significantly smaller compared to the original CSN dataset used for training other attention-based models for code understanding, such as CodeBERT.
StarCoder Dataset. The StarCoder Dataset (Li et al., 2023) is a cleaned and de-duplicated dataset with Personally Identifiable Information (PII) redacted. The dataset contains git-issues and git-commits along with 86 programming languages. Additionally, the authors add additional information to every sample such as GitHub star, repository name and path. We remove these information along with the accompanying tokens such as <gh-stars> before using the dataset. To create the subset used in our work, we simply select the first $N$ samples, where $N=300k$ or $600k$.
Finetuning Datasets. The details about finetuning datasets is presented in Table 6.
All pretraining and finetuning data used in the paper are publicly available.
Table 6: Size of Finetuning datasets
Benchmarks Train Validation Test
NLCodeSearch 251K 9.6K 19K
Clone Detection 900K 416K 416K
Type Inference 9M 201K 224K
SQA 251K 9.6K 19K
Devign 21K 2.7K 2.7K
DiverseVul 419K 52K 52K
Complexity Prediction 3613 452 452
Appendix D Other Training Setup and Details
In this section, we briefly mention different training setups we experimented with and their impact on performance.
Learning Rate (LR) scheduler. We experimented with cosine and linear LR schedulers for both pretraining and finetuning. The CodeSSM performed the best with cosine scheduler in pretraining and linear scheduler in finetuning
Warm-up. Warm-up is typically used with transformers to prevent over-fitting to data seen in the early stages of training. We first experimented with no warm-up and a warm-up step of of 1% of total training steps. We found that CodeSSM performed better on MLM validation with no warm-up. Subsequently, we experiment with a fixed warm-up step of 300 and surprisingly, with this small warm-up step, the pretraining performance improved. Further experiments are required to understand how warm-up impacts SSM training.
Learning rate We experimented with learning rates in proximity to what is used for transformers on code intelligence tasks. In general, SSMs performed better with lower learning rates compared to those for transformers. We used the learning rate of $5e^{-5}$ for pretraining for all the parameters (including SSM parameters).
Weight Decay The weight decay for SSMs is set to 0. The weight decay for Bias and Normalization layer is also set to 0. For the rest weight decay is set to 0.01.
Larger batch sizes. We used gradient accumulation to increase the batch size. However, with larger batch sizes, the CodeSSM model performed worse. This could be due to less number of gradient updates with larger batch sizes and it’s possible that larger batch sizes help with larger datasets.
Tokenizer We use the CodeT5p-220M tokenizer Wang et al. (2023b) from transformers library  [^7] https://huggingface.co/docs/transformers/v4.17.0/en/index, which has a vocabulary size of 32100, for two reasons. Since it is designed specifically for code, the CodeT5 tokenizer does not significantly increase the input context length after tokenization. Moreover, it has been used for the CodeSearchNet dataset, which we use for pretraining.
Weight Initialization The layers of the CodeSSM model are initialized using the corresponding layers of the BiGS model. BiGS is a 23 layer model. For CodeSSM with 12 layers, we discard the weights of additional layers of BiGS. We initialize the embedding layer of CodeSSM randomly because its vocabulary differs from that of BiGS. The task-specific heads are also initialized randomly.
Appendix E Pretraining evaluation
In Table 7, we present the pretraining accuracy of BertCoder and multiple CodeSSM variants on a hold-out set from the CodeSearchNet dataset.
Table 7: MLM accuracy of different models, trained with a context window of 2048, on validation set from CSN dataset.
Model Accuracy
BertCoder 84.230
CodeSSM 80.022
CodeSSM-do 79.180
CodeSSM-pos 79.640
CodeF 73.324
Appendix F Theoretical basis for improvements in performance
SSMs outperform Transformers due to two characteristics of their architectural design:
Figure 4: Magnitude Plot of Fourier Transform of SSM kernel for 12 layers of CodeSSM
Figure 5: Phase Plot of Fourier Transform of SSM Kernel for 12 layers of CodeSSM
  • The initialization of A matrix Theoretically, the initialization with HiPPO matrix allows SSMs to include unlimited historical information from a sequence. Additionally, with bi-directionality, the CodeSSM is also able to encode future information with no limit. Practically, the unlimited historical information (or context) means the input length. As we show in our paper, the pretraining length does not limit this ability and SSMs can extrapolate to longer lengths after pretraining.
  • Transfer function of SSMs Given that SSMs are Linear Time-Invariant (LTI) systems, we can look at their transfer function to understand how the input affects the output. We can write this relation as:
    $\mathbf{Y}=\mathbf{H(\omega)}*\mathbf{U(\omega)}$ (7)
    where $\mathbf{Y}$ represents the output, $\mathbf{H(\omega)}$ and $\mathbf{U(\omega)}$ represents the Fourier transforms of the SSM kernel and the input respectively. Since, Fourier Transforms are complex,
    $\begin{split}\mathbf{H(\omega)}=\mathbf{M}\angle\theta\\ \mathbf{M}\in[0,\infty]\\ \theta\in[0,180]\end{split}$ (8)
    We can write the self-attention equation in a similar way,
    $\mathbf{Y}=\mathbf{A(u)}*\mathbf{T(u)}$ (9)
    where $\mathbf{A(u)}=Softmax(QK^{T}/\sqrt{d})$ and $\mathbf{T(u)}\in[0,1]$ which is some transformation of the input.
    In SSMs, a phase shift of 180 degree between kernel and input results in forgetting some information (when added to the input through skip connection), while unbounded magnitude allows SSMs to weigh some important information very strongly. On the other hand, the transformers cannot forget any input information even if those are not relevant for a task.
The plots of magnitude and phase of all layers of the pretrained CodeSSM model for a kernel of length 10 is shown in Figure 4 and Figure 5 respectively. The plot shows forward kernel on the left and backward kernel on the right.
Do note, however, that $\mathbf{A(u)}$ explicitly calculates relation between each pair of input tokens, while $\mathbf{H(\omega)}$ does not. This results in a key limitation of SSMs which is that they fail to encode complex local relations. This failure can result in slight degradation in performance on tasks requiring local information with complex relations among tokens, as evidenced by performance on Type Inference (Table 2).
let itemStatus={intervalId:null,
status:"Uploading",
urn:"",
derivativeUrn:""};
item.formData=itemStatus;
let derivativeUrn=jobStatus1
.derivatives
.filter(derivative)
=>derivative
.outputType$=$"ifc")[0]
.children[0].urn;
Appendix G Qualitative Analysis of Failure cases in Type Inference
We investigated the failure cases of CodeSSM on the type inference task and found that CodeSSM often fails at predicting the type when the local context is very complicated. For example, consider the code line in the listing above.
Here the type of derivativeUrn is string but CodeSSM predicted it as Output while BertCoder predicted it correctly. In the same code however, CodeSSM predicted the correct type for string variables when the local context was simpler. Incorrect prediction of string type in complex contexts is also the reason for lower Top 100 F1.
To predict the correct type, the model must look at the location from where the value of the variable is coming, i.e, the model should encode which tokens influence the value of a given token. Each SSM kernel is a mixture of multiple wavelengths. The influence of one token on another depends on the frequency coefficients at a specific wavelength Hua et al. (2025b). Since our SSM implementation has only one kernel per SSM block, there are only few wavelengths with high coefficients. In Figure 4, it can be observed that the later layers has higher coefficients for lower frequency (longer wavelengths) and hence, focus on long range dependencies. Since, type inference require both long and short range dependency, the model fails at encoding sufficient information. Including multiple kernels in SSM blocks can potentially alleviate this issue.
Appendix H Additional Results
Here are some additional results on Type Inference (Table 8) with Top 100 F1 and on Clone Detection (Table 9) with Precision and Recall.
Table 8: Results on type inference in terms of overall F1 and top-100 F1. The best performance is in bold and other noteworthy results are underlined.
Model Context Length Overall F1 Top 100 F1
CodeBERT 512 0.595 0.850
BertCoder 2048 0.666 0.816
RoCoder 256 0.687 84.17
CodeSSM-base 256 0.588 0.769
CodeSSM-base 512 0.603 0.772
CodeSSM-base 2048 0.615 0.792
CodeSSM-pos 2048 0.594 0.766
CodeSSM-do 2048 0.609 0.771
CodeSSM-Sc 2048 0.619 0.780
CodeSSM-Sc-large 2048 0.637 0.800
Table 9: Results on code clone detection in terms of Precision, Recall and F1 score. The best performance is in bold and other noteworthy results are underlined.
Model Context Length Precision Recall F1
CodeBERT 512 0.947 0.934 0.941
BertCoder 2048 0.928 0.958 0.942
RoCoder 256 0.934 0.944 0.938
CodeSSM-base 256 0.888 0.929 0.908
CodeSSM-base 512 0.949 0.942 0.946
CodeSSM-base 2048 0.940 0.926 0.935
CodeSSM-pos 2048 0.921 0.959 0.939
CodeSSM-do 2048 0.926 0.937 0.931
CodeSSM-Sc 2048 0.950 0.926 0.938
CodeSSM-Sc-large 2048 0.933 0.942 0.940

### v3

@@
 Abstract
-Although transformers are widely used for various code-specific tasks, they have some significant limitations.
-In this paper, we investigate State Space Models (SSMs) as a potential alternative to transformers for code understanding tasks, such as code retrieval, classification, and clone detection.
-Previous research has already demonstrated that SSMs are more compute-efficient than transformers.
-In our work, we show that SSMs are also more sample-efficient and can effectively extrapolate to longer contexts (beyond the pretraining context) during fine-tuning.
-Through comprehensive experiments, we demonstrate that SSMs could serve as a viable alternative to transformers for code understanding tasks, while addressing some of the major limitations associated with transformers.
+Although transformers dominate many code-specific tasks, they have significant limitations.
+This paper explores State Space Models (SSMs) as a promising alternative for code understanding tasks such as retrieval, classification, and clone detection.
+We introduce CodeSSM, the first SSM-based model trained on code corpora to assess its effectiveness.
+Our results demonstrate that SSMs are more sample-efficient and can extrapolate to longer contexts beyond the pretraining length.
+Extensive experiments show that SSMs offer a viable alternative to transformers, addressing several their limitations.
+Additionally, CodeSSM reduces memory usage by up to 64% compared to transformers at a context length of 2048, with greater savings as context length grows.
 CodeSSM: Towards State Space Models for Code Understanding
@@
 1 Introduction
-Transformers (Vaswani et al., 2017) have been known to perform well on various code intelligence tasks.
-One of the main reasons for this improvement is the pretraining - finetuning paradigm used to train transformer models, which is possible due to the ease of parallelization of attention-based transformer architectures.
-Under this paradigm, the transformer model is initially trained on a large corpus of unlabeled data using a self-supervised training objective.
-This is followed by a supervised finetuning process on a smaller labeled dataset.
+Transformers (Vaswani et al., 2017) have been known to perform well in various applications.
+One of the main reasons for this improvement is the pretraining - fine-tuning paradigm used to train transformer models, which is possible due to the ease of parallelization of attention-based transformer architectures.
+Under this paradigm, the transformer model is initially trained on a large corpus of unlabeled data using a self-supervised training objective, followed by a supervised fine-tuning process on a smaller labeled dataset.
 However, the performance gain offered by the transformer models comes with some trade-offs, such as quadratic complexity, a substantial data requirement, and high inference costs.
-While methods such as linear attention (Katharopoulos et al., 2020) and sparse attention (Guo et al., 2022b; Zaheer et al., 2020; Condevaux and Harispe, 2022) have been proposed to address computational inefficiency, the practical gains from these methods are limited (Yang et al., 2025; Qin et al., 2022).
+Although methods such as linear attention (Katharopoulos et al., 2020) and sparse attention (Guo et al., 2022b; Zaheer et al., 2020; Condevaux and Harispe, 2022) have been proposed to address computational inefficiency, the practical gains from these methods are limited (Yang et al., 2025; Qin et al., 2022).
 Another limitation of transformers is their fixed context window resulting from positional embedding.
@@
 One-dimensional Convolutional Neural Networks (1D CNNs) are a fast and positional embedding-free alternative to transformers.
-1D CNNs, when trained using the pretraining-finetuning approach, show improved performance on some tasks but fail to capture long-range dependencies in input data (Tay et al., 2021).
+1D CNNs, when trained using the pretraining-finetuning approach, show improved performance on some tasks, but do not capture long-range dependencies in input data (Tay et al., 2021).
+Figure 1: Memory consumption and throughput of CodeSSM and RoCoder model during inference.
 State-space models (SSMs)Gu et al.
@@
 SSMs leverage the strengths of 1D CNNs while effectively capturing long-range dependencies with linear-time complexity.
-Additionally, SSMs are position-aware, eliminating the need for positional embedding.
+In addition, SSMs are position-aware, eliminating the need for positional embedding.
 In this work, we explore the application of SSMs to the domain of code understanding.
@@
 We also investigate various training aspects, such as the effects of positional embedding, dropout, bidirectional SSMs, and varying pretraining context lengths.
-We compare the CodeSSM with two transformer models that we train – (1) BertCoder: based on Bert architecture (Devlin et al., 2019) and (2) RoCoder: based on RoFormer architecture Su et al.
+We compare CodeSSM with two transformer models that we train – (1) BertCoder: based on the Bert architecture (Devlin et al., 2019) and (2) RoCoder: based on the RoFormer architecture Su et al.
 (2024).
-We assess the performance of our model against several high-performing transformers as well.
-Our study demonstrates that SSMs exhibit strong sample efficiency in code language modeling tasks.
-Notably, CodeSSM achieves over 50% accuracy in masked language modeling (MLM) with fewer than 3,000 training steps.
-Additionally, the code language modeling capabilities of CodeSSM also transfer well to downstream tasks as CodeSSM outperforms transformers on multiple code understanding tasks when pretrained under similar conditions.
+Moreover, we compare our model with Zamba, a hybrid SSM transformer model based on the Zamba2 architecture Glorioso et al.
+(2024).
+We also assess the performance of our model against several high-performing transformers, which are trained with larger data and complex pretraining objectives.
+Our study demonstrates that SSMs exhibit strong sample efficiency in code modeling tasks.
+In particular, CodeSSM achieves more than 50% precision in Masked Language Modeling (MLM) with fewer than 3,000 training steps.
+Furthermore, we show that CodeSSM consumes up to 64% less memory compared to transformers and is significantly faster (see Figure˜1).
+The efficiency in terms of speed and memory consumption in CodeSSM (compared to transformers) improves with increasing input length.
+Additionally, the code language modeling capabilities of CodeSSM also transfer well to downstream tasks, as CodeSSM outperforms transformers on multiple code understanding tasks when pretrained under similar conditions.
 CodeSSM is also competitive with (and even surpasses) transformer models trained on significantly larger datasets and with more complex pretraining objectives.
 Furthermore, another advantage of SSMs is their ability to extrapolate to much larger contexts than those encountered during pretraining, effectively addressing a key limitation of the transformer architecture.
-Our objective in training the CodeSSM model is not to achieve state-of-the-art performance on coding benchmarks, but to explore and document the strengths and limitations of SSMs in understanding code.
-This goal also motivates our usage of a small dataset and MLM pretraining which allows us to quickly experiment with different model configurations and compare them under similar training settings.
-Although we observe improvements in performance with larger pretraining data, we will leave the exploration of very large training datasets and more complex training objectives for future work.
+Our goal in training the CodeSSM model is not to achieve state-of-the-art results on coding benchmarks, but to explore and document the strengths and limitations of SSMs for code understanding.
+This motivated our use of a small dataset and MLM pretraining, enabling rapid experimentation with and comparison of different model configurations under consistent training conditions.
+Although we observed that larger pretraining datasets improved performance, the investigation of extensive datasets and more complex training objectives is left for future work.
 Summary of contributions:
-  • We conduct the first systematic evaluation of State Space Models (SSMs) for code understanding, introducing a specialized model, which we term CodeSSM.
-Our analysis reveals that CodeSSM achieves robust performance across multiple benchmarks, consistently outperforming attention-based models while requiring substantially less pretraining data.
-  • Our study presents the first empirical evidence that State Space Models (SSMs) achieve superior sample efficiency in code-understanding tasks compared to transformers.
+  • We present the first systematic evaluation of State-Space Models (SSMs) for code understanding, introducing a specialized model, which we term CodeSSM.
+Our analysis reveals that CodeSSM achieves robust performance on retrieval, classification and clone detection tasks, consistently outperforming attention-based models while requiring substantially less pretraining data and memory.
+  • Our study presents the first empirical evidence that SSMs achieve superior sample efficiency in code-understanding tasks compared to transformers.
 CodeSSM substantially outperforms transformer baselines in data-constrained scenarios with only a few thousand training samples.
@@
 CodeSSM also performs very well on the adversarial text-code search task.
-Together these observations indicate that SSMs exhibit robust semantic understanding capabilities.
+Together, these observations indicate that SSMs exhibit robust semantic understanding capabilities.
+The code is available here.
 2 CodeSSM
@@
 CodeSSM-dropout.
-Dropout layers (Srivastava et al., 2014) prevent over-fitting in neural networks.
-So, to investigate over-fitting in CodeSSM, we create a variant with two dropout masks, positioned after each multiplication operations (see Figure 1).
+Dropout layers (Srivastava et al., 2014) prevent overfitting in neural networks.
+So, to investigate overfitting in CodeSSM, we create a variant with two dropout masks, positioned after each multiplication operation (see Figure˜2).
 We refer to this variant as CodeSSM-do.
 CodeF and CodeSSM-Uni.
-Recent research Lee-Thorp et al.
-(2022) questioned the necessity of self-attention in transformers.
-To investigate the significance of SSM, we conducted a similar ablation study by replacing SSM blocks in CodeSSM with DFT (see Figure 1), creating a variant we call CodeF.
+Recent research questioned the necessity of self-attention in transformers Lee-Thorp et al.
+(2022).
+To investigate the significance of SSM, we conducted a similar ablation study by replacing SSM blocks in CodeSSM with DFT (see Figure˜2), creating a variant we call CodeF.
 Furthermore, we analyze the importance of bidirectional processing via CodeSSM-Uni (Unidirectional Gated SSM), a variant of CodeSSM that removes flip operations from the CodeSSM layer.
@@
 Dataset.
-We initially trained CodeSSM and a BERT-like transformer model, which we refer to as BertCoder, on 1.8M text-code pairs from the CodeSearchNet (CSN) dataset.
+We initially trained CodeSSM and a BERT-like transformer model, which we refer to as BertCoder, on 1.8M text code pairs from the CodeSearchNet (CSN) dataset.
 The small pretraining dataset allows us to quickly experiment with various model configurations.
-Both models were trained with a large context window of 2048, enabling them to be fine-tuned on benchmarks with lengths shorter than 2048.
+Both models were trained with a context window of 2048, enabling them to be fine-tuned on benchmarks with lengths shorter than 2048.
 We refer to the CodeSSM model trained with the CSN dataset as CodeSSM-base.
-Additionally, we also train CodeSSM variants such as CodeSSM-pos, CodeSSM-do, CodeF and CodeSSM-Uni (see 2.1) on the same dataset and context length.
-Context Length and Length Extrapolation.
-To investigate the length extrapolation capability of SSMs, we pretrain the CodeSSM model with a 256-length context window.
+We also train CodeSSM variants such as CodeSSM-pos, CodeSSM-do, CodeF and CodeSSM-Uni (see 2.1) on the same dataset and context length.
+Context Length and Length Extrapolation To investigate the length extrapolation capability of SSMs, we pre-train CodeSSM with a context window of 256 length.
 Since we cannot change the context window for BertCoder once pretrained with a smaller context, we pretrained a RoFormer model Su et al.
-(2024), which is similar to the BERT model but with Rotatory Positional Embeddings (RoPE), with a 256-length context window.
+(2024), which is similar to the BERT model but with Rotatory Positional Embeddings (RoPE), with a context window of length 256.
 We refer to this model as RoCoder.
-CSN dataset separates docstring and code, which creates an issue when they are concatenated and subsequently truncated.
+The CSN dataset separates the docstring and code, which creates an issue when they are concatenated and subsequently truncated.
 This can result in one modality (text) being much longer than the other modality (code).
-We also observed this discrepancy results in reduction in performance in tasks involving two sequences (refer to column 3 of Table 2).
-To alleviate this issue, we use a subset of 1.8M samples from the StarCoder dataset (Li et al., 2023) to train the models with 256-length context (referred to as CodeSSM-sc-256).
+We also observed that this discrepancy results in a reduction in performance in tasks involving two sequences (refer to column 3 of Table˜2).
+To alleviate this issue, we use a subset of 1.8M samples from the StarCoder dataset (Li et al., 2023) to train models with 256-length context (referred to as CodeSSM-sc-256).
 This subset contains 300k samples from the six programming languages (PLs) in the CSN dataset  [^2] Go, Java, Javascript, PHP, Python, Ruby.
@@
 Data Scaling.
-To study the impact of large datasets on the downstream performance of CodeSSM, we also train it on a large subset of the StarCoder dataset with 4.8M samples (referred as CodeSSM-sc-L).
+To study the impact of large datasets on the downstream performance of CodeSSM, we also train it in a large subset of the StarCoder dataset with 4.8M samples (referred to as CodeSSM-sc-L-2048).
 This subset contains 600k samples from 8 PLs (C and C# along with CSN PLs).
@@
 3 Experiments
-In this section, we describe the benchmarks on which we evaluate CodeSSM and the baselines with which we compare our model.
+In this section, we describe baselines to which we compare CodeSSM and the benchmarks on which we evaluate it.
 3.1 Baselines
 We trained two transformer-based models as baselines: a large context (2048) BERT model to compare SSMs’ large context abilities and a small context RoFormer model to compare the length extrapolation abilities of transformers and SSMs.
-The models being compared are trained under the same settings: same dataset, same objective, context length, and batch size.
+We also train a Zamba model to compare our model with the other SSM variants.
+The models are trained in the same settings: the same data set, the same objective, the same context length, and batch size.
 Additionally, we also mention the performance of some well-performing transformer models in each task to place CodeSSM in the context of SOTA encoder-only transformer models.
-However, it should be noted that these transformer models are trained on a much larger dataset and with a mixture of multiple training objectives (see Table 1).
+However, it should be noted that these transformer models are trained on a much larger dataset and with a mixture of multiple training objectives (see Table˜1).
 3.2 Benchmarks
@@
 This task is difficult because the model has to locate the correct code throughout the test set.
-In SQA, a StackOverflow question is given and the goal is to retrieve the highest upvoted answer.
+In SQA, a StackOverflow question is given, and the goal is to retrieve the highest upvoted answer.
 This dataset is derived from the original StackOverflow dataset  [^3] https://www.kaggle.com/datasets/stackoverflow/stacksample/data.
-The average length of question and answer are 1.2k and 1.4k respectively.
+The average lengths of question and answer are 1.2k and 1.4k, respectively.
 Sequence classification tasks.
@@
 (2023).
-Vulnerability detection is a binary classification task where the goal is to identify whether a given code contains security vulnerabilities.
+Vulnerability detection is a binary classification task in which the goal is to identify whether a given code contains security vulnerabilities.
 The Devign benchmark Zhou et al.
@@
 Token classification task.
-In the type-inference Jesse and Devanbu (2022) task, the model has to predict the type (such as integer, boolean, etc.) of each token in a given code.
+In the type-inference task Jesse and Devanbu (2022), the model has to predict the type (such as integer, boolean, etc.) of each token in a given code.
 The label consists of 50K most frequent (built-in and user-defined) types, while the remaining types have been replaced with the UNK type.
@@
 All models are fine-tuned using a context length of 512 for text-to-code search, clone detection, type inference, and DiverseVul.
-For the Devign benchmark, the models are trained with a context length of 400.
+For the Devign benchmark, the models are fine-tuned with a context length of 400.
 In SQA and complexity prediction, we use a context length of 1024.
@@
 RoCoder 256 29.44 0.938 0.687
+Zamba 256 - 0.232 0.503
 StarCoder CodeSSM-Sc 256 28.41 0.948 0.624
@@
 CodeSSM-Sc-L 2048 30.87 0.940 0.637
-Table 3: Results on SQA and APPS benchmark in terms of MRR. The best performance is in bold and other noteworthy results are underlined.
+Table 3: Results on SQA benchmark in terms of MRR. The best performance is in bold and other noteworthy results are underlined.
 Model SQA
@@
 RoCoder 80.91
+Zamba -
 CodeSSM-Sc-256 82.12
@@
 4 Results
-We present the performance of our model on various benchmarks described in Section 3.2
+We present the performance of our model on various benchmarks described in Section˜3.2
 4.1 Retrieval Tasks
 NLCodeSearch(AdvTest).
-Column 3 of Table 2 presents results of various models evaluated on NLCodeSearch benchmark.
+Column 3 of Table˜2 presents the results of various models evaluated in NLCodeSearch benchmark.
 Among the models trained on the CSN dataset, CodeSSM-base (with a context size of 2048) outperforms other models, including CodeBERT, which was pretrained with significantly more data.
-CodeSSM-base with a context length of 256 performs poorly compared to other variants – we explain the reason for this in Section 2.2.
+CodeSSM-base with a context length of 256 performs poorly compared to other variants – we explain the reason for this in Section˜2.2.
 CodeSSM-sc-256 shows remarkable performance and length extrapolation to double the length on which it was trained and is competitive with RoCoder and CodeSSM-sc-2048.
+However, Zamba – the other SSM based model – failed to train on retrieval task.
+We attribute the inability to train on retrieval-based tasks to the causal convolution in the Mamba block, which prevents bidirectional understanding.
 These results are particularly noteworthy given the adversarial nature of the NLCodeSearch benchmark, where critical semantic identifiers such as function and variable names are obfuscated with generic placeholder tokens.
@@
 StackOverflowQA.
-Table 3 presents the performance of CodeSSM in conjunction with the baseline models trained by us, BertCoder and RoCoder, on the SQA benchmark.
+Table˜3 presents the performance of CodeSSM in conjunction with the baseline models trained by us on the SQA benchmark – BertCoder and RoCoder.
 SQA allows us to evaluate the models with longer contexts.
-CodeSSM-sc with a pretraining context length of 256 achieves strong results – it successfully extrapolates to contexts four times longer than those seen during pretraining.
+CodeSSM-sc with a pretraining context length of 256 achieves strong results: it successfully extrapolates to contexts four times longer than those seen during pretraining.
 Furthermore, CodeSSM-sc outperforms BertCoder (pretrained with a context length of 2048) by 6 points, despite being pretrained with a context length eight times shorter.
@@
 Furthermore, we observe that increasing the pretraining context length leads to further improvements in CodeSSM’s performance.
+Similarly to NLCodeSearch, Zamba did not train on SQA as well.
 Table 4: Results on DiverseVul benchmark in terms of F1 macro. The best performance is in bold and other noteworthy results are underlined.
@@
 RoCoder 68.15 53.47
+Zamba 59.43 56.55
 CodeSSM-Sc-256 69.31 64.53
@@
 RoCoder 93.06
+Zamba 90.16
 CodeSSM-do-2048 92.82
@@
 Devign.
-Column 3 of Table 4 shows the results of the various models on the devign benchmark.
+Column 3 of Table˜4 shows the results of the various models on the devign benchmark.
 CodeSSM outperforms BertCoder and RoCoder as well as CodeBERT, GraphCodeBERT, and CodeT5.
@@
 However, in the case of CodeSSM, the smaller context model performs much better.
-Notably, RoCoder performs poorly, potentially due to over-fitting, as the devign dataset is very small (only 21k training samples).
+Notably, RoCoder performs poorly, potentially due to overfitting, as the devign dataset is very small (only 21k training samples).
 RoPE, used in RoCoder, is also prone to positional overfitting Liu et al.
@@
 (2025).
+Despite being a unidirectional model, Zamba performs slightly better than RoCoder.
 DiverseVul.
-Column 2 of Table 4 shows that CodeSSM and its variants outperform BertCoder and RoCoder on the DiverseVul benchmark, which is a larger vulnerability detection dataset compared to Devign.
+Column 2 of Table˜4 shows that CodeSSM and its variants outperform BertCoder and RoCoder on the DiverseVul benchmark, which is a larger vulnerability detection dataset compared to Devign.
 It shows that CodeSSM can perform better than transformers even on a larger dataset.
-CodeSSM even outperforms CodeBERT, GraphCodeBERT, and CodeT5, which are trained on a much larger dataset and have multiple pretraining objectives (see Table 1).
+CodeSSM even outperforms CodeBERT, GraphCodeBERT, and CodeT5, which are trained on a much larger dataset and have multiple pretraining objectives (see Table˜1).
 Complexity Prediction.
-Table 5 shows that CodeSSM-sc-256 outperforms all the transformer models even the ones trained with longer context length (BertCoder) and larger data and more pretraining objectives (CodeBERT, GraphCodeBERT, UnixCoder).
+Table˜5 shows that CodeSSM-sc-256 outperforms all transformer models, even those trained with a longer context length (BertCoder) and larger data and more pretraining objectives (CodeBERT, GraphCodeBERT, UnixCoder).
 Additionally, CodeSSM-sc-2048 shows very strong performance as pretraining on longer context length improves the performance further on this task.
 4.3 Token classification task (Type inference)
-Column 5 of Table 2 shows the results on the type inference benchmark.
+Column 5 of Table˜2 shows the results on the type inference benchmark.
 CodeSSM (across all context lengths) matches or exceeds CodeBERT.
 However, both RoCoder and BertCoder outperform CodeSSM.
-We provide reasons for poor performance of CodeSSM on type inference in Section 5.5 and Appendix G.
+We provide reasons for the poor performance of CodeSSM in type inference in Section˜5.6 and Appendix˜G.
 4.4 Clone Detection
-Column 4 of Table 2) shows the F1 on clone detection task.
+Column 4 of Table˜2 shows the F1 in the clone detection task.
 The CodeSSM-sc with a pretraining context length of 256 outperforms both CodeBERT, BertCoder, and RoCoder in terms of F1.
-Under a similar training setup (2048 context on CSN), removing positional embedding does have a minimal impact on performance, but with the advantage that the maximum input length need not be fixed during pretraining.
-The decrease in performance after introducing dropout shows the model does not overfit despite the dataset being unbalanced.
+Under a similar training setup (2048 context in CSN), removing positional embedding does have a minimal impact on performance, but with the advantage that the maximum input length need not be fixed during pretraining.
+The decrease in performance after introducing dropout shows that the model does not overfit despite the data set being unbalanced.
 5 Research Questions
@@
 Figure 3: MLM training accuracy of CodeSSM and BertCoder, trained with a context window of 2048. The CodeSSM model achieves decent MLM accuracy in very few training steps.
-Figure 2 shows the MLM training accuracy trajectories of CodeSSM and BertCoder, both trained with a maximum context of 2048.
+Figure˜3 shows the MLM training accuracy trajectories of CodeSSM and BertCoder, both trained with a maximum context of 2048.
 CodeSSM demonstrates rapid initial learning, achieving more than 50% accuracy in only 3,000 training steps (96,000 training samples).
-BertCoder, on the other hand, exhibits better long-term learning characteristics, outperforming CodeSSM by 4% points after six training epochs (see Appendix E).
-However, CodeSSM outperforms transformers in all downstream tasks except type inference, as discussed in Section 4.
-The performance on downstream tasks shows that the code modeling capability observed during pretraining transfers well to multiple downstream tasks.
+BertCoder, on the other hand, exhibits better long-term learning characteristics, outperforming CodeSSM by 4% points after six training epochs (see Appendix˜E).
+However, CodeSSM outperforms transformers in all downstream tasks except type inference, as discussed in Section˜4.
+Performance in downstream tasks shows that the code modeling capability observed during pretraining transfers well to multiple downstream tasks.
 We evaluated CodeSSM on two small benchmarks: The Devign vulnerability detection benchmark with only 21k training samples and the complexity prediction benchmark with only 3.6k training samples.
@@
 (2023a).
-We also present some (informal) theoretical basis for CodeSSM’s performance and the magnitude and phase spectrum of Fourier Transform of SSM kernel in Appendix F.
+We also present some (informal) theoretical basis for CodeSSM’s performance and the magnitude and phase spectrum of Fourier transform of SSM kernel in Appendix˜F.
 5.2 How well does CodeSSM extrapolate to longer contexts?
-We studied the transformer models and CodeSSMs for positional embedding and different pretraining context lengths to investigate their length extrapolation capabilities.
+We studied transformer models and CodeSSMs for positional embedding and different pretraining context lengths to investigate their length extrapolation capabilities.
 On multiple tasks, we observe that BertCoder outperforms CodeBERT despite less pretraining data simply due to the longer pretraining context length.
@@
 Moreover, with a longer pretraining context, CodeSSM outperforms RoCoder on NLCodeSearch as well.
-Therefore, with more computational resources, CodeSSM can be efficiently trained with longer contexts to enhance the performance on retrieval tasks.
-However, it’s important to note that longer pretraining contexts do not benefit all downstream tasks; in fact, performance slightly decreases on vulnerability detection and clone detection.
+Therefore, with more computational resources, CodeSSM can be efficiently trained with longer contexts to improve the performance on retrieval tasks.
+However, it is important to note that longer pretraining contexts do not benefit all downstream tasks; in fact, performance slightly decreases in vulnerability detection and clone detection.
 Furthermore, we observe that CodeSSM performs better without positional embedding.
 CodeSSMs without positional embeddings do not face any length limitation, unlike transformers with absolute positional embedding, or positional biases observed with different methods such as Alibi or RoPE (Liu et al., 2024b; Goel et al., 2025).
+5.3 How fast and memory efficient is CodeSSM?
+Table 6: Memory Consumption during Inference (in GB) on various tasks with varying context lengths.
+Model NLCodeSearch (512) SQA (1024) SQA (2048)
+CodeSSM 9.33 15.84 28.84
+RoCoder 9.87 25.36 79.12
+The usage of GPU memory during inference is shown in Table˜6.
+CodeSSM model consumes 5.5% less memory than the transformer model (RoCoder) on the NLCodeSearch task where the context length is 512.
+The difference in memory consumption between CodeSSM and RoCoder increases to 37.5% on SQA task where the context length is 1024.
+As the input length increases to 2048, CodeSSM saves 63.55% memory compared to the transformers.
+These findings are in line with previous works on SSM Le Bronnec et al.
+(2024).
+Table 7: Inference throughput (samples per second) with varying context lengths.
+Model NLCodeSearch (512) SQA (1024) SQA (2048)
+CodeSSM 36.38 22.50 11.49
+RoCoder 38.11 19.76 8.48
+Transformers are slightly faster compared to SSM in small input sizes (<=512) but SSMs are faster as the input length increases.
 5.4 Can SSMs be replaced by DFT?
 We ask this question because Lee-Thorp et al.
-(2022) showed that self-attention can be replaced with DFT in transformers with minimal reduction on NLP tasks.
+(2022) showed that self-attention can be replaced with DFT in transformers with minimal reduction in NLP tasks.
 We also observed that in all tasks RoCoder performed significantly better than BertCoder.
 Hua et al.
-(2025b) showed that RoPE also performs token-level (non-uniform) DFT.
+(2025b) showed that RoPE also performs token-level (nonuniform) DFT.
 Additionally, SSMs also implicitly perform DFT of input and kernel during convolution operation.
 To answer this question, we performed an ablation study by replacing the SSM transformation in CodeSSM with DFT and creating a unidirectional variant of CodeSSM.
-The results are presented in Figure 3 (and in Table 7).
+The results are presented in Figure˜4 (and in Table˜10).
 We observe that both variants perform significantly worse compared to CodeSSM.
@@
 First, we study the impact of adding dropout to the CodeSSM layer.
-Across all tasks, including the complexity prediction task which has only 3.6k samples, the performance of the model reduces when dropout is added.
-The performance degradation suggests that CodeSSM does not overfit during training regardless of the size of the finetuning dataset.
+Across all tasks, including the complexity prediction task, which has only 3.6k samples, the performance of the model decreases when dropout is added.
+This performance degradation suggests that CodeSSM does not overfit during training regardless of the size of the finetuning dataset.
 Second, we evaluated the model on the adversarial NLCodeSearch task, by which informative tokens such as function and variable names have been replaced with noninformative generic tokens.
@@
 We discuss two possible scenarios where the performance of SSMs on code understanding tasks can be inferior to that of the transformers.
-As discussed in Section 5.1, transformers can outperform SSMs as the number of samples increases.
+As discussed in Section˜5.1, transformers can outperform SSMs as the number of samples increases.
 This phenomenon can hinder the performance of SSMs on some tasks such as Type Inference.
@@
 However, when the number of samples is large and the context length of finetuning datasets is long, SSMs can surpass transformers in such scenarios (as seen in the SQA benchmark).
-SSMs can effectively capture long-range dependencies but do not model local dependencies (Zuo et al., 2024) very well.
-The type inference benchmark requires the model to understand both short and long-range dependencies.
+SSMs can effectively capture long-range dependencies, but they do not model local dependencies (Zuo et al., 2024) very well.
+The type inference benchmark requires the model to understand both short- and long-range dependencies.
 For example, in the code sample x = y+1; z = func(x), the type of x can be inferred from the local context (variable y), while the type of z can only be inferred from the return type of the function func, which requires capturing long-range dependencies.
 We also observe that this problem becomes more challenging when the local context is complicated.
-We provide a qualitative example of CodeSSM’s failure in the task of type inference in Appendix G.
-We can potentially improve CodeSSM’s performance on this task by making the step size of SSMs (see Appendix A) data dependent Qi et al.
-(2024) or with multiple SSM kernels (Appendix G), which can improve its understanding of short-range dependencies.
+We provide a qualitative example of the failure of CodeSSM in the type inference task in Appendix˜G.
+We can potentially improve CodeSSM’s performance on this task by making the step size of SSMs (see Appendix˜A) data dependent Qi et al.
+(2024) or with multiple SSM kernels (Appendix˜G), which can improve its understanding of short-range dependencies.
+5.7 Expressivity gaps in SSMs and their fixes
+Some recent theoretical works document the expressivity limitations of SSMs.
+In particular, Merrill et al.
+(2024) shows that SSMs cannot express some important problems that require state tracking.
+This class of problems includes evaluating Python code.
+However, it should be noted that transformers have also been shown to not express this class of problems while a single layer RNN can Merrill and Sabharwal (2023).
+However, the expressive power of SSMs can be improved through different architectural improvements.
+For example, Merrill et al.
+(2024) showed that making the input of the transition matrix independent or adding nonlinearities can extend the expressive power of SSM.
+Similarly, Nishikawa and Suzuki (2025) showed that when SSMs are preceded and followed by feed-forward layers, they have the same dynamic token abilities as transformers.
+Another limitation is the requirement of specific reparameterization.
+Agarwal et al.
+(2024) proposed a new SSM formulation by augmenting spectral filters with negative eigenvalues, which removed the requirement of reparameterization.
+The SSM variant used in CodeSSM is S4D.
+S4D also has negative eigenvalues but still requires specific reparameterization.
+The best way to parametrize SSM and their practical benefits remain an open research question Wang and Li (2024).
 6 Related Work
@@
 Dao et al.
-(2022; 2024) introduced hardware-level optimizations, but the benefits of these optimizations are hardware-specific and do not necessarily apply to newer or different hardware architectures Shah et al.
+(2022) and Dao (2024) introduced hardware-level optimizations, but the benefits of these optimizations are hardware-specific and do not necessarily apply to newer or different hardware architectures Shah et al.
 (2024).
@@
 (2024) and Wang et al.
-(2024) enhances the base transformer models for longer contexts using repository-level information retrieval, while Clement et al.
+(2024) improve base transformer models for longer contexts using repository-level information retrieval, while Clement et al.
 (2021); Cheng et al.
@@
 (2024a); Ma et al.
-(2024) use static analysis tools.
-The usage of static analysis tools is complementary to CodeSSM, as the performance of CodeSSM can also be further improved through the use of these tools.
+(2024) use static analysis tools for repository-level tasks.
+The usage of static analysis tools is complementary to CodeSSM, as the performance of CodeSSM can also be further improved by using these tools.
 Additionally, CodeSSM can be a drop-in replacement as an efficient retriever model.
@@
 In this work, we investigated the effectiveness of SSM in code understanding tasks by introducing CodeSSM.
-Our results show significant potential of CodeSSM in code modeling and understanding.
+Our results show a significant potential for CodeSSM in code modeling and understanding.
 Our findings show that SSMs can be a robust and sample-efficient alternative to transformers, offering the added benefit of length extrapolation.
-While the training pipeline and architectural improvements for transformers have evolved extensively over the past decade, the optimal training conditions for SSMs have yet to be thoroughly investigated.
+Although the training pipeline and architectural improvements for transformers have evolved extensively in the past decade, the optimal training conditions for SSMs have yet to be thoroughly investigated.
 This work represents a step toward understanding these conditions and uncovering the advantages of SSMs.
-Notably, the performance gains achieved by CodeSSM were obtained under suboptimal conditions, indicating the potential for further enhancements.
+In particular, the performance gains achieved by CodeSSM were obtained under suboptimal conditions, indicating the potential for further enhancements.
 With advancements in training methodologies and architectural designs for SSM-based models, their performance could be significantly improved.
 Limitations
-While we have done an extensive study of SSMs for code understanding, our work has some limitations.
+Although we have done an extensive study of SSMs for code understanding, our work has some limitations.
   • Single kernel transformation: We have used a simple implementation of S4d Gu et al.
 (2022a) in the CodeSSM layer.
-This implementation has only one SSM kernel shared across all input channel.
-Incorporating multiple kernels in an SSM block can further improve the performance.
-  • Short-range dependencies: We have discussed SSMs’ limitations in understanding short-range dependencies with respect to type inference task.
-However, we do not implement a solution to this problem.
-Some methods to alleviate this issue have been proposed in literature.
-We leave the exploration of these methods in the context of code understanding for future work.
+This implementation has only one SSM kernel shared across all input channels.
+Incorporating multiple kernels into an SSM block can further improve performance.
+  • Short-range dependencies: We have discussed SSMs’ limitations in understanding short-range dependencies in the context of the type inference task.
+However, we do not provide a solution to this problem.
+While some methods have been proposed to alleviate this issue in the literature, we leave the their exploration in the context of code understanding for future work.
   • Generative tasks: CodeSSM currently uses SSMs only in convolution mode.
@@
 CodeSearchNet dataset can contain personally identifiable information (PII).
-While StarCoder dataset attempts to remove PII, there might still be some PII in the dataset.
-It is not clear whether our model can leak such data if present in the pre-training dataset, but we acknowledge that such potential risk exist.
-Further, our work involved training multiple models.
-While the model and dataset sizes were small, the training can still have environmental impacts.
+While the StarCoder dataset attempts to remove PII, there might still be some PII in the dataset.
+It is not clear whether our model can leak such data if present in the pre-training dataset, but we acknowledge that such potential risk exists.
+In addition, our work involved training multiple models.
+Although the size of the model and the data set was small, training can still have environmental impacts.
 Appendix A Background
@@
 (2023a) is a bidirectional gated architecture which consists of two SSM layers as shown in Fig.
-1.
+2.
 This gated architecture is adapted from the gated attention unit described in Hua et al.
@@
 To create the subset used in our work, we simply select the first $N$ samples, where $N=300k$ or $600k$.
-Finetuning Datasets. The details about finetuning datasets is presented in Table 6.
+Finetuning Datasets. The details about finetuning datasets is presented in Table˜8.
 All pretraining and finetuning data used in the paper are publicly available.
@@
 In this section, we briefly mention different training setups we experimented with and their impact on performance.
+Table 9: Hyperparameters for CodeSSM, BertCoder, RoCoder and Zamba models
+Hyperparameters CodeSSM BertCoder RoCoder Zamba
+Number of Layers 12 12 12 14
+Hidden Dimension 1024 1024 1032 1024
+Positional Embedding No Yes Yes No
+Pretraining context length 256 2048 256 256
 Learning Rate (LR) scheduler.
@@
 Appendix E Pretraining evaluation
-In Table 7, we present the pretraining accuracy of BertCoder and multiple CodeSSM variants on a hold-out set from the CodeSearchNet dataset.
+In Table˜10, we present the pretraining accuracy of BertCoder and multiple CodeSSM variants on a hold-out set from the CodeSearchNet dataset.
 Table 10: MLM accuracy of different models, trained with a context window of 2048, on validation set from CSN dataset.
@@
 On the other hand, the transformers cannot forget any input information even if those are not relevant for a task.
-The plots of magnitude and phase of all layers of the pretrained CodeSSM model for a kernel of length 10 is shown in Figure 4 and Figure 5 respectively.
+The plots of magnitude and phase of all layers of the pretrained CodeSSM model for a kernel of length 10 is shown in Figure˜5 and Figure˜6 respectively.
 The plot shows forward kernel on the left and backward kernel on the right.
@@
 This results in a key limitation of SSMs which is that they fail to encode complex local relations.
-This failure can result in slight degradation in performance on tasks requiring local information with complex relations among tokens, as evidenced by performance on Type Inference (Table 2).
+This failure can result in slight degradation in performance on tasks requiring local information with complex relations among tokens, as evidenced by performance on Type Inference (Table˜2).
 let itemStatus={intervalId:null,
@@
 =>derivative
-.outputType$=$"ifc")[0]
+.outputType==="ifc")[0]
 .children[0].urn;
@@
 Since our SSM implementation has only one kernel per SSM block, there are only few wavelengths with high coefficients.
-In Figure 4, it can be observed that the later layers has higher coefficients for lower frequency (longer wavelengths) and hence, focus on long range dependencies.
+In Figure˜5, it can be observed that the later layers has higher coefficients for lower frequency (longer wavelengths) and hence, focus on long range dependencies.
 Since, type inference require both long and short range dependency, the model fails at encoding sufficient information.
@@
 Appendix H Additional Results
-Here are some additional results on Type Inference (Table 8) with Top 100 F1 and on Clone Detection (Table 9) with Precision and Recall.
+Here are some additional results on Type Inference (Table˜11) with Top 100 F1 and on Clone Detection (Table˜12) with Precision and Recall.
 Table 11: Results on type inference in terms of overall F1 and top-100 F1. The best performance is in bold and other noteworthy results are underlined.